OpenStudy (anonymous):

Andy is waiting at the signal. As soon as the light turns green, he accelerates his car at a uniform rate of 8.00 meters/second2. What is the displacement of Andy’s car after 30.0 seconds?

OpenStudy (anonymous):

X=(at^2)/2 X=(8.30^2)/2 X=3600 meters.